# Improved subsample-and-aggregate via the private modified winsorized mean

### v1

@@ Appendix D Zero-concentrated DP quantiles (zCDP)
Consider neighboring datasets $\mathbf{x}_{n},\mathbf{x}_{n}^{\prime}$ with empirical distribution functions, $F_{n}$ and $F_{n}^{\prime}$, respectively. Let ${\rm D}_{\alpha}(P||Q)$ denotes the $\alpha$-Renyi Divergence between two measures $P$ and $Q$. That is,
@@
To this end, let $K$ be the random integer at which the process stops on the original dataset $\mathbf{x}_{n}$, and let $K^{\prime}$ be the random integer at which the process stops on the neighboring dataset $\mathbf{x}_{n}^{\prime}$. It follows by post-processing, (see Lemma 15 of [8]) that ${\rm D}_{\alpha}(\tilde{\xi}_{q,\mathbf{x}_{n}}||\tilde{\xi}_{q,\mathbf{x}_{n}^{\prime}})\leq{\rm D}_{\alpha}(K||K^{\prime})$, since $\tilde{\xi}_{q,\mathbf{x}_{n}^{\prime}}=\beta^{K}+\ell-1$. It suffices to show that ${\rm D}_{\alpha}(K||K^{\prime})\leq\rho\alpha$ for all $\alpha>1$.
To this end, let $M\stackrel{{\scriptstyle\text{d}}}{{=}}K$, let $\nu_{j}$ be the joint distribution of the sequence of random variables $Z=(K,V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$ and let $\nu_{j}^{\prime}$ be the distribution of $Z^{\prime}=(K^{\prime},V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$. Take $I$ to be the function returning the first element, so that $I(Z)=K$ and $I(Z^{\prime})=K^{\prime}$. It follows by Lemma 15 of [8] (post-processing) that
@@
We now consider ${\rm D}_{\alpha}(\nu_{j}||\nu_{j}^{\prime})$ for arbitrary $j$. Let $Q_{j}$ and $Q_{j}^{\prime}$ denote the distribution of $K$ and $K^{\prime}$ conditional on $(V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$. Let $\mu_{j}$ be the joint distribution of the sequence of random variables $(V_{1}/n\sqrt{\rho},\ldots,V_{j-1},V_{j+1}/n\sqrt{\rho},\ldots)$. Next, we have that by Lemma 15 of [8] (composition),
@@
It then suffices to show that, for any $j\geq 0$, we have that ${\rm D}_{\alpha}(Q_{j}||Q_{j}^{\prime})\leq\alpha\rho$. To this end, let $(v_{1}/n\sqrt{\rho},\ldots,v_{j-1}/n\sqrt{\rho},v_{j+1}/n\sqrt{\rho},\ldots)$ be the realizations of $(V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$. Next, for any positive integer $k$, let $z_{k}=F_{n}(\beta^{k}+\ell-1)$, $\tau_{k}=\max_{i<k}[F_{n}(\beta^{i}+\ell-1)+v_{i}]$, $z_{k}^{\prime}=F^{\prime}_{n}(\beta^{k}+\ell-1)$ and $\tau_{k}^{\prime}=\max_{i<k}[F_{n}^{\prime}(\beta^{i}+\ell-1)+v_{i}]$. To prove the result, we first have that
@@
×$\hskip 200.0003pt\times\varphi(n\sqrt{\rho}(u_{1}+\tau_{k}-z_{k}+z_{k}^{\prime}-\tau_{k}^{\prime}))du_{1}du_{2}$
@@
$\Pr\left(K^{\prime}=k\Big{|}\{V_{i}=v_{i}\}_{i=1}^{k-1},\{V_{i}=v_{i}\}_{i=k+1}^{\infty}\right)$ $=\Pr\left(\tau_{k}^{\prime}\leq\hat{q}<z_{k}^{\prime}+V_{k}/n\sqrt{\rho}\Big{|}\{V_{i}=v_{i}\}_{i=1}^{k-1},\{V_{i}=v_{i}\}_{i=k+1}^{\infty}\right)$
$=h_{k}((n\sqrt{\rho})^{-1}(V-V_{j})).$
@@
where the second last line uses the fact that the Renyi divergence between two normal distributions with variance $\sigma^{2}$ and means $\mu_{1},\mu_{2}$ is given by $\frac{1}{2}\alpha(\mu_{1}-\mu_{2})^{2}/2\sigma^{2}.$ Now, we just need to bound the expectation $\mathbb{E}(U_{3})$, from which it follows by definition that $-1/n<\mathbb{E}(U_{3})<1/n$. As a result, we have that ${\rm D}_{\alpha}(Q_{j}||Q_{j}^{\prime})\leq\alpha\rho,$ which completes the proof. ∎
@@ Appendix E Proving finite sample bounds for private quantiles
$1\leq\beta\leq\frac{u-\hat{\xi}_{q-t,X}+1}{u-\hat{\xi}_{q+t,X}+1}\iff 1\leq\beta\leq\frac{-\ell+\hat{\xi}_{1-(q-t),Y}+1}{-\ell+\hat{\xi}_{1-(q+t),Y}+1}\iff 1\leq\beta\leq\frac{\hat{\xi}_{q^{\prime}+t,Y}-\ell+1}{\hat{\xi}_{q^{\prime}-t,Y}-\ell+1}.$
@@
$\Pr\left(|q-F_{n}(\tilde{\xi}_{q})|>t\right)$ $\leq\left(\frac{\log(\hat{\xi}_{q-t}+1-\ell)}{\log\beta}+1\right)\exp\left(-n(t\vee t^{2})(\sqrt{\rho_{1}\vee\rho_{1}^{2}}\wedge\sqrt{\rho_{2}\vee\rho_{2}^{2}})/4\right),$
@@
$\Pr\left(|q-F_{n}(\tilde{\xi}_{q})|>t\right)$ $\leq\left(\frac{\log(u-\hat{\xi}_{q+t}+1)}{\log\beta}+1\right)\exp\left(-n(t\vee t^{2})(\sqrt{\rho_{1}\vee\rho_{1}^{2}}\wedge\sqrt{\rho_{2}\vee\rho_{2}^{2}})/4\right).$
Proof of Lemma E.2.
The proof mirrors that of Lemma E.1. We start by demonstrating the first inequality. Let $\hat{q}^{\prime}=q+V/n\sqrt{\rho}_{1}$. Observe that
@@
$=\Pr\left(V_{i}/n\sqrt{\rho_{2}}+F_{n}((\hat{\xi}_{q+t}-\ell+1)+\ell-1)\leq q+V/n\sqrt{\rho_{1}}\right)$
@@
Then, for the second inequality, the same argument as in Lemma E.1 applies. ∎
Appendix F Error in the proof of Theorem 1 of Lugosi and Mendelson, [28]
Here, we demonstrate that there is an error in the proof of Theorem 1 of Lugosi and Mendelson, [28]. Namely, one step uses the inequalities
$\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2\tilde{X}}}(X)\leq\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\geq\xi_{1-\zeta/2,X}\right\})\text{ and }\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2,X}}(X)\geq-\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\leq\xi_{2\zeta,X}\right\}),$
which do not hold, see page 400 and 401 of Lugosi and Mendelson, [28]. We show they do not hold by counterexample. For the remainder of this section, we use the notation of Lugosi and Mendelson, [28].
F.1 The lower bound
@@
$\hskip 100.00015pt+\mathbb{E}(\overline{X}(1-\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})}-\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})}))$
@@
$\hskip 100.00015pt-\mathbb{E}(\overline{X}\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})})-\mathbb{E}(\overline{X}\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})})$
@@
$\hskip 100.00015pt+\mathbb{E}((Q_{1-\epsilon/2}(\overline{X})-\overline{X})\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})}).$
@@ F.2 The upper bound
$\mathbb{E}\phi_{\mu+Q_{2\epsilon}(\overline{X}),\mu+Q_{1-\epsilon/2}(\overline{X})}(X)\leq\mu+\mathbb{E}[\overline{X}\mathbbm{1}_{\overline{X}>Q_{1-\epsilon/2}(\overline{X})}],$ (F.1)
@@
$\mathbb{E}\phi_{\mu+Q_{2\epsilon}(\overline{X}),\mu+Q_{1-\epsilon/2}(\overline{X})}(\overline{X})$ $=(p_{1}+p_{2})x_{2}+(p_{3}+p_{4})x_{3}$
@@
To prove that (F.1) does not hold for $F$, we must show that $((p_{1}+p_{2})x_{2}+(p_{3}+p_{4})x_{3})>x_{4}p_{4}$ for appropriate $p_{i}$ and $x_{i}$. So we just need a solution to the set of constraints:
$2\displaystyle 0<\epsilon<1/$ 2
$\epsilon<2(p_{4}+p_{3})$
$p_{1}<2\epsilon<p_{2}+p_{1}$
$x_{1}<x_{2}<0<x_{3}<x_{4}$
$p^{\prime}1_{4}=1$
$x^{\prime}p=0$
$1\displaystyle 0<p_{i}<$ 1
$((p_{1}+p_{2})x_{2}+(p_{3}+p_{4})x_{3})>x_{4}p_{4}.$
@@
Appendix G Skewed distributions and trimmed means

### v3

@@ Appendix D Zero-concentrated DP quantiles (zCDP)
Consider neighboring datasets $\mathbf{x}_{n},\mathbf{x}_{n}^{\prime}$ with empirical distribution functions, $F_{n}$ and $F_{n}^{\prime}$, respectively. Let ${\rm D}_{\alpha}(P||Q)$ denotes the $\alpha$-Rényi Divergence between two measures $P$ and $Q$. That is,
@@
To this end, let $K$ be the random integer at which the process stops on the original dataset $\mathbf{x}_{n}$, and let $K^{\prime}$ be the random integer at which the process stops on the neighboring dataset $\mathbf{x}_{n}^{\prime}$. It follows by post-processing, (see Lemma 15 of Bun and Steinke, (2016)) that ${\rm D}_{\alpha}(\tilde{\xi}_{q,\mathbf{x}_{n}}||\tilde{\xi}_{q,\mathbf{x}_{n}^{\prime}})\leq{\rm D}_{\alpha}(K||K^{\prime})$, since $\tilde{\xi}_{q,\mathbf{x}_{n}^{\prime}}=\beta^{K}+\ell-1$. It suffices to show that ${\rm D}_{\alpha}(K||K^{\prime})\leq\rho\alpha$ for all $\alpha>1$.
To this end, let $M\stackrel{{\scriptstyle\text{d}}}{{=}}K$, let $\nu_{j}$ be the joint distribution of the sequence of random variables $Z=(K,V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$ and let $\nu_{j}^{\prime}$ be the distribution of $Z^{\prime}=(K^{\prime},V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$. Take $I$ to be the function returning the first element, so that $I(Z)=K$ and $I(Z^{\prime})=K^{\prime}$. It follows by Lemma 15 of Bun and Steinke, (2016) (post-processing) that
@@
We now consider ${\rm D}_{\alpha}(\nu_{j}||\nu_{j}^{\prime})$ for arbitrary $j$. Let $Q_{j}$ and $Q_{j}^{\prime}$ denote the distribution of $K$ and $K^{\prime}$ conditional on $(V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$. Now, let $\mu_{j}$ be the joint distribution of the sequence of random variables $(V_{1}/n\sqrt{\rho},\ldots,V_{j-1},V_{j+1}/n\sqrt{\rho},\ldots)$. Next, we have that by Lemma 15 of Bun and Steinke, (2016) (composition),
@@
It then suffices to show that, for any $j\geq 0$, we have that ${\rm D}_{\alpha}(Q_{j}||Q_{j}^{\prime})\leq\alpha\rho$. To this end, let $(v_{1}/n\sqrt{\rho},\ldots,v_{j-1}/n\sqrt{\rho},v_{j+1}/n\sqrt{\rho},\ldots)$ be the realizations of the sequence of random variables $(V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$. Next, for any positive integer $k$, let $z_{k}=F_{n}(\beta^{k}+\ell-1)$, $\tau_{k}=\max_{i<k}[F_{n}(\beta^{i}+\ell-1)+v_{i}]$, $z_{k}^{\prime}=F^{\prime}_{n}(\beta^{k}+\ell-1)$ and $\tau_{k}^{\prime}=\max_{i<k}[F_{n}^{\prime}(\beta^{i}+\ell-1)+v_{i}]$. To prove the result, we first have that
@@
×$\hskip 100.00015pt\times\varphi(n\sqrt{\rho}(u_{1}+\tau_{k}-z_{k}+z_{k}^{\prime}-\tau_{k}^{\prime}))du_{1}du_{2}$
@@
$\Pr\left(K^{\prime}=k\Big{|}\{V_{i}=v_{i}\}_{i=1}^{k-1},\{V_{i}=v_{i}\}_{i=k+1}^{\infty}\right)$
$\hskip 60.00009pt=\Pr\left(\tau_{k}^{\prime}\leq\hat{q}<z_{k}^{\prime}+V_{k}/n\sqrt{\rho}\Big{|}\{V_{i}=v_{i}\}_{i=1}^{k-1},\{V_{i}=v_{i}\}_{i=k+1}^{\infty}\right)$
$\hskip 60.00009pt=h_{k}((n\sqrt{\rho})^{-1}(V-V_{j})).$
@@
where the second last line uses the fact that the Rényi divergence between two normal distributions with variance $\sigma^{2}$ and means $\mu_{1},\mu_{2}$ is given by $\frac{1}{2}\alpha(\mu_{1}-\mu_{2})^{2}/2\sigma^{2}.$ Now, we just need to bound the expectation $\mathbb{E}(U_{3})$, from which it follows by definition that $-1/n<\mathbb{E}(U_{3})<1/n$. As a result, we have that ${\rm D}_{\alpha}(Q_{j}||Q_{j}^{\prime})\leq\alpha\rho,$ which completes the proof. ∎
@@ Appendix E Proving finite sample bounds for private quantiles
$1\leq\beta\leq\frac{u-\hat{\xi}_{q-t,X}+1}{u-\hat{\xi}_{q+t,X}+1}\iff 1\leq\beta\leq\frac{-\ell+\hat{\xi}_{1-(q-t),Y}+1}{-\ell+\hat{\xi}_{1-(q+t),Y}+1}\\ \iff 1\leq\beta\leq\frac{\hat{\xi}_{q^{\prime}+t,Y}-\ell+1}{\hat{\xi}_{q^{\prime}-t,Y}-\ell+1}.$
@@
$\Pr\left(|q-F_{n}(\tilde{\xi}_{q})|>t\right)$ $\leq\left(\frac{\log(\hat{\xi}_{q-t}+1-\ell)}{\log\beta}+1\right)$
×$\hskip 20.00003pt\times\exp\left(-n(t\vee t^{2})(\sqrt{\rho_{1}\vee\rho_{1}^{2}}\wedge\sqrt{\rho_{2}\vee\rho_{2}^{2}})/4\right),$
@@
$\Pr\left(|q-F_{n}(\tilde{\xi}_{q})|>t\right)$ $\leq\left(\frac{\log(u-\hat{\xi}_{q+t}+1)}{\log\beta}+1\right)$
×$\hskip 20.00003pt\times\exp\left(-n(t\vee t^{2})(\sqrt{\rho_{1}\vee\rho_{1}^{2}}\wedge\sqrt{\rho_{2}\vee\rho_{2}^{2}})/4\right).$
Proof of Lemma 3.
The proof mirrors that of Lemma 2. We start by demonstrating the first inequality. Let $\hat{q}^{\prime}=q+V/n\sqrt{\rho}_{1}$. Observe that
@@
$=\Pr\left(\frac{V_{i}}{n\sqrt{\rho_{2}}}+F_{n}((\hat{\xi}_{q+t}-\ell+1)+\ell-1)\leq q+\frac{V}{n\sqrt{\rho_{1}}}\right)$
@@
Then, for the second inequality, the same argument as in Lemma 2 applies and the proof is complete. ∎
Appendix F Proof of Corollary 1
Proof of Corollary 1.
First, the triangle inequality implies
$|\tilde{T}(\mathbf{X}_{n}^{\prime})^{\prime}-T(F)|<|\mathrm{Bias}_{n,\eta}(T)|+|\tilde{T}(\mathbf{X}_{n}^{\prime})^{\prime}-\mathbb{E}(T(\mathbf{X}_{n}^{\prime(1)}))|.$
Next, given that the conditions are satisfied by assumption, we apply Theorem 1. ∎
Appendix G Error in Lugosi and Mendelson (2021)
Here, we demonstrate that there is an error in the proof of Theorem 1 of Lugosi and Mendelson, (2021). Namely, one step uses the inequalities
$\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2\tilde{X}}}(X)\leq\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\geq\xi_{1-\zeta/2,X}\right\}),$
and
$\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2,X}}(X)\geq-\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\leq\xi_{2\zeta,X}\right\}),$
which do not hold, see page 400 and 401 of Lugosi and Mendelson, (2021). We show they do not hold by counterexample. For the remainder of this section, we use the notation of Lugosi and Mendelson, (2021).
G.1 The lower bound
@@
$\hskip 27.00005pt+\mathbb{E}(\overline{X}(1-\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})}-\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})}))$
@@
$\hskip 27.00005pt-\mathbb{E}(\overline{X}\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})})-\mathbb{E}(\overline{X}\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})})$
@@
$\hskip 27.00005pt+\mathbb{E}((Q_{1-\epsilon/2}(\overline{X})-\overline{X})\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})}).$
@@ G.2 The upper bound
$\mathbb{E}\phi_{\mu+Q_{2\epsilon}(\overline{X}),\mu+Q_{1-\epsilon/2}(\overline{X})}(X)\leq\mu+\mathbb{E}[\overline{X}\mathbbm{1}_{\overline{X}>Q_{1-\epsilon/2}(\overline{X})}],$ (10)
@@
$\mathbb{E}\phi_{\mu+Q_{2\epsilon}(\overline{X}),\mu+Q_{1-\epsilon/2}(\overline{X})}(\overline{X})=(p_{1}+p_{2})x_{2}+(p_{3}+p_{4})x_{3},$
and, also,
@@
To prove that (10) does not hold for $F$, we must show that $((p_{1}+p_{2})x_{2}+(p_{3}+p_{4})x_{3})>x_{4}p_{4}$ for appropriate $p_{i}$ and $x_{i}$. So we just need a solution to the set of constraints given by:
$0<\epsilon<1/2,\ \epsilon<2(p_{4}+p_{3}),\ p_{1}<2\epsilon<p_{2}+p_{1},\ x_{1}<x_{2}<0<x_{3}<x_{4},$
[…]
@@
Appendix H Skewed distributions and trimmed means